# De-Hallucinator: Iterative Grounding for LLM-Based Code Completion

### v1

@@ 5.1.2. Dataset and Ground Truth
We construct a dataset of API-related code completion tasks by removing API usages from the benchmark projects and by considering the removed code as the ground truth to be predicted by a model. Specifically, we use LibCST [^9] https://libcst.readthedocs.io/en/latest/ to identify API calls that use an API from another file in the same project. For each such API call, we remove the lines containing the call. If a call spans multiple lines, we remove all of them. To prevent data leakage from imports of the API in the ground truth, we also remove all local imports (i.e. starting with .) and imports of the package itself. Next, we check if the off-the-shelf LLMs can predict the exact code as in the original file using the code preceding the cursor as the prompt. If an LLM predicts exactly the original code, we ignore this API usage for the evaluation, as there is no need to further improve the prediction. We continue with this process for each of the four models, until we have ten code completion tasks for each of the eleven projects. During this process, we ignore 18, 51, 76, and 31 completions for UniXcoder, CodeGen, CodeGen v2.5, and StarCoder+, respectively. Moreover, for StarCoder+ 26 completions result in non-ascii completions, which we also ignore. Overall, the evaluation dataset consists of 11 projects $\times$ 10 $\times$ 4 models = 440 code completion tasks.
During the static pre-analysis, we ensure to not index any information that would not be available during the code completion task. This is necessary because we run our CodeQL queries (Section 3.3) on a database that contains all of the code in the project, including the function that we want to complete. To avoid any leakage of information, the static pre-analysis does not extract any information from function bodies, except for __init__ functions, which are analyzed to extract attribute references, and which we therefore exclude from the evaluation dataset.
@@ 5.1.3. Metrics
  • Edit distance. To quantify the number of edits a developer would have to apply after receiving a code completion, we measure the edit distance between the predicted code and the ground truth. This metric provides a sense for how many token edits are saved when using De-Hallucinator. We compute edit distance using the Levenshtein distance at the subtoken level. For each pair of completion and ground truth, we tokenize the code pieces with a GPT-2 fast tokenizer, [^10] https://huggingface.co/docs/transformers/model_doc/gpt2#transformers.GPT2TokenizerFast and then calculate the edit distance using NLTK’s edit_distance. [^11] https://www.nltk.org/_modules/nltk/metrics/distance.html#edit_distance That is, given predicted code $c_{pred}$, a ground truth $c_{correct}$, and the tokenizer function $tokens$, the metric is:
    $edit\_distance(tokens(c_{pred}),tokens(c_{correct}))$
  • Normalized edit similarity. Similar to previous work (Lu et al., 2022) we also compute the normalized edit similarity. To this end, we normalize the absolute edit distance (computed as above) to the length of the longer of the two token sequences, and then turn the result into a similarity metric:
    $1-\frac{edit\_distance(tokens(c_{pred}),tokens(c_{correct}))}{max(|tokens(c_{pred})|,|tokens(c_{correct})|)}$
  • Exact API match. Since the goal of De-Hallucinator is to predict better API usages, we measure how many of all desired API usages are predicted exactly as in the ground truth. To identify the API usages in the lines of code to complete, we extract function calls, including the access path to the function, and the parameters. For example, given a line of code
    docs = ds.find_by_keyword(keyword)}
[…]
    python
ds.find_by_keyword(keyword). The exact API match then is the percentage of exact matches between the prediction and the ground truth API usages. Let $api(c)$ be the API usages in code $c$, then the metric is:
$\frac{|api(c_{pred})\cap api(c_{correct})|}{|api(c_{correct})|}$
@@ 5.1.4. Hardware
We perform the experiments with the CodeGen v2.5 model on a machine equipped with two Nvidia T4 GPUs, each having 16GB of memory. The experiments with the UniXcoder, CodeGen, and the StarCoder+ models are performed on a machine with a single Nvidia Tesla V100 with 32GB of memory. Each machine has a 48-core Intel Xeon CPU clocked at 2.20GHz.
@@ 5.2. RQ1: Effectiveness of De-Hallucinator
Table 3. Effectiveness of De-Hallucinator compared to the baseline on four off-the-shelf LLMs. The bold numbers show statistically significant improvement over the baseline (Type 1). The numbers in parentheses show the relative improvement over the baseline (Type 1).
Metric
Prompt type
UniXCoder (125M)
CodeGen v1 (2B)
CodeGen v2.5 (7B)
StarCoder+ (15B)
Edit distance (lower is better)
Type 1
52.39
40.04
47.21
44.60
Type 2
46.80 (10.67%)
33.42 (16.53%)
31.64 (32.98%)
35.94 (19.42%)
Type 3
25.86 (50.64%)
30.72 (23.28%)
30.09 (36.26%)
33.50 (24.89%)
Normalized edit similarity (higher is better)
Type 1
33.40
43.64
43.93
33.24
Type 2
37.31 (11.71%)
48.00 (9.99%)
49.40 (12.45%)
38.00 (14.32%)
Type 3
42.58 (27.48%)
48.93 (12.12%)
50.15 (14.16%)
39.66 (19.31%)
Exact API match (higher is better)
Type 1
4.77
7.12
8.33
5.68
Type 2
4.77 (0.0%)
10.15 (42.56%)
11.59 (39.13%)
7.50 (32.04%)
Type 3
5.91 (23.90%)
11.06 (55.34%)
13.41 (60.98%)
7.50 (32.04%)
@@
In the first set of experiments, we investigate to what extent De-Hallucinator improves code completions compared to the baseline. By default, we run De-Hallucinator with $k=4$ iterations and add $n=20$ API references into the prompt. As shown in 3, De-Hallucinator reduces the edit distance by 9.32 to 26.53 tokens, on average, which is between 23.28% and 50.64% relative improvement. This in turn translates to normalized edit similarity improvements of 12.12% to 27.48% relative to the baseline. Moreover, the approach relatively improves the exact API matches by 23.90% to 60.98%. For example, for the CodeGen v2.5 model, De-Hallucinator is able to predict 1.5 times more APIs correctly than the baseline. We evaluate the statistical significance of our improvements using the Wilcoxon test, with the Pratt method to deal with zero values. De-Hallucinator shows statistically significant improvements over the baseline consistently for all metrics and all models. The fact that the highest improvement is achieved for the exact API match metric shows that De-Hallucinator effectively achieves its goal of improving API usages in code completions. Given the importance of this problem in realistic code completion scenarios (Section 2), De-Hallucinator has the potential to significantly improve the productivity of developers.
5.2.1. Examples
7 shows the initial completion of CodeGen, with only the preceding code as prompt. In this example the model uses the assert statement instead of the custom assert_equivalent function, which is defined in the project. After augmenting the prompt with the API reference of this function, the same model correctly predicts the call to the existing API, as shown in 7.
def test_idempotent_any_syntatically_valid_python(
src_contents: str, mode: black.FileMode
) -> None:
dst_contents = black.format_str(src_contents, mode=mode)
# And check that we got equivalent and stable output.
[…]
# API Reference:
# check_stability_and_equivalence( src_contents: str, dst_contents: str, *, mode: Mode) -> None # Perform stability and equivalence checks. Raise AssertionError if source and destination contents ar
# assert_equivalent(src: str, dst: str) -> None # Raise AssertionError if src and dst arent equivalent.
%**** main.tex Line 675 ****…
def test_idempotent_any_syntatically_valid_python(
src_contents: str, mode: black.FileMode
) -> None:
dst_contents = black.format_str(src_contents, mode=mode)
# And check that we got equivalent and stable output.
[…]
Figure 7. Completion by CodeGen (baseline) highlighted in red, the ground truth highlighted in green, and the completion by De-Hallucinator using the augmented prompt highlighted in blue.
8 shows another scenario where De-Hallucinator improves the completion. In this case the correct function is used by the model in the first try, but the order of parameters is wrong. By providing the API reference in the prompt, De-Hallucinator predicts the correct API usage, as shown in 8.
async def schedule_formatting(sources: Set[Path], fast: bool, write_back: WriteBack,
mode: Mode, report: "Report", loop: asyncio.AbstractEventLoop,
executor: "Executor") -> None:
"""Run formatting of sources in parallel using the provided executor‘.
%**** main.tex Line 700 ****(Use ProcessPoolExecutors for actual parallelism.)
write_back‘, fast‘, and mode options are passed to
:func:‘format_file_in_place‘.
cache: Cache = {}
if write_back not in (WriteBack.DIFF, WriteBack.COLOR_DIFF):
cache = read_cache(mode)
[…]
# API Reference:
# filter_cached(cache: Cache, sources: Iterable[Path]) -> Tuple[Set[Path], Set[Path]] # Split an iterable of paths in sources into two sets. The first contains paths of files that modifi
async def schedule_formatting(sources: Set[Path], fast: bool, write_back: WriteBack,
mode: Mode, report: "Report", loop: asyncio.AbstractEventLoop,
executor: "Executor") -> None:
"""Run formatting of sources in parallel using the provided executor‘.
(Use ProcessPoolExecutors for actual parallelism.)
write_back‘, fast‘, and mode options are passed to
:func:‘format_file_in_place‘.
cache: Cache = {}
if write_back not in (WriteBack.DIFF, WriteBack.COLOR_DIFF):
cache = read_cache(mode)
[…]
@@ 5.3. RQ2: Correct Retrieval of API References
To better understand the effectiveness of De-Hallucinator, we investigate how often the approach successfully augments the prompt with the correct API references. Answering this question for all completion tasks and all LLMs is difficult, because comparing the API references to the API usages is non-trivial due to different ways of importing APIs and passing arguments. Instead, we manually inspect a sample of 20 completion tasks per LLM and count the number of times an API used in the ground truth is successfully added to the prompt by De-Hallucinator.
@@
Missing/wrong Expected API ref. added
in init. comp. Type 2 Type 3 (k=2) Type 3 (k=3) Type 3 (k=4)
@@
In our inspected samples, there are in total between 22 and 25 API usages (a completion can contain multiple API usages). The new prompt generated by our approach contains the correct API between two and six times, as shown in 4. Moreover, for CodeGen and CodeGen v2.5 there are five completion tasks where the completion from prompt of type 1 either misses the APIs or uses them incorrectly, but in the completions from type 2 or type 3 prompts, the API reference section of the prompt contains the correct API. The same happens for UniXcoder in four tasks and for StarCoder+ in two tasks. Overall, the results show that De-Hallucinator is able to successfully augment the prompt with the correct API references in many cases. For cases where the approach fails to add the correct API reference into the prompt, the main reason is that the initial completion has low relevance w.r.t. the ground truth.
@@ 5.4. RQ3: Impact of the Hyperparameters
This research question evaluates the impact of the two main parameters of our approach. First, we consider the number $k$ of iterations of the main loop of De-Hallucinator, where we run the approach with $k\in\{2,3,4,5\}$. As shown in 9, the first iteration provides significant improvement on all metrics, but the gain is reduced with further iterations. Higher values of $k$ are beneficial when the model cannot immediately predict a relevant completion, but upon presenting the first round of API references, the model responds with a more relevant completion. At the same time, even $k=2$ provides clear improvements over the baseline, which makes De-Hallucinator useful even in scenarios where the cost of querying the model is high. We choose $k=4$ for all other experiments in this paper, since $k=5$ does not provide large enough improvements on our metrics.
Figure 9. Relative improvements over the baseline for the maximum number of iterations, $k$.
Figure 10. Relative improvements of De-Hallucinator for different maximum API reference counts, $n$.
Second, we study the impact of the maximum number $n$ of API references that we add to the prompt. Setting low values for $n$ can result in missing relevant context, whereas adding many API references reduces the space left for context before the cursor. We perform experiments with $n\in\{2,10,20,40\}$. 10 shows the results on our three metrics, with the similarity metrics increasing with $n$, and exact API match peaking between $n=10$ and $n=20$. As a default in the rest of the paper, we use $n=20$.

### v3

@@ 5.1.3. Datasets
With the goal of having a diverse set of projects in terms of size, domain, and popularity, we gather a dataset of eleven public Python projects from GitHub, shown in Table 2 for the code completion task. We construct a dataset of API-related code completion tasks by removing API usages from the benchmark projects and by considering the removed code as the ground truth to be predicted by a model. For each such API call, we remove the lines containing the call. If a call spans multiple lines, we remove all of them. To prevent data leakage from imports of the API in the ground truth, we also remove API-related imports. Next, we check if the off-the-shelf LLMs can predict the exact code as in the original file using the code preceding the cursor as the prompt. If an LLM predicts exactly the original code, we ignore this API usage for the evaluation, as there is no need to further improve the prediction and to avoid any potential memorizations. We continue with this process for each of the four models, until we have ten code completion tasks for each of the eleven projects. During this process, we ignore 18, 51, 76, and 31 completions for UniXcoder, CodeGen, CodeGen v2.5, and StarCoder+, respectively. Overall, the code completion evaluation dataset consists of 11 projects $\times$ 10 $\times$ 4 models = 440 code completion tasks.
Test Generation
For the test generation task, we also gather a diverse set of 12 JavaScript projects from GitHub, shown in Table 2. These projects cover a variety of domains, such as website generation, code formatting, and process management. Since TestPilot cannot generate tests for ES modules, we only consider JavaScript projects that are CommonJS packages. The rest of the setup is the same as in the TestPilot paper (Schäfer et al., 2024).
@@ 5.1.4. Metrics
  • Edit distance. To quantify the number of edits a developer needs to apply after receiving a code completion, we measure the edit distance between the predicted code and the ground truth. This metric provides a sense for how many token edits are saved when using De-Hallucinator. We compute edit distance using the Levenshtein distance at the subtoken level. For each pair of completion and ground truth, we tokenize the code pieces with a GPT-2 fast tokenizer, [^9] https://huggingface.co/docs/transformers/model_doc/gpt2#transformers.GPT2TokenizerFast and then calculate the edit distance using NLTK’s edit_distance. [^10] https://www.nltk.org/_modules/nltk/metrics/distance.html#edit_distance
  • Normalized edit similarity. Similar to previous work (Lu et al., 2022) we also compute the normalized edit similarity. To this end, we normalize the absolute edit distance (computed as above) to the length of the longer of the two token sequences, and then turn the result into a similarity metric.
  • Exact API match. Since the goal of De-Hallucinator is to predict better API usages, we measure how many of all desired API usages are predicted exactly as in the ground truth. To identify the API usages in the lines of code to complete, we extract function calls, including the access path to the function, and the parameters.
@@
Test Generation
For the test generation task, we use the following three metrics:
  • Number of passing tests. We count the number of passing tests as a proxy for code without any hallucinations.
  • Coverage. We measure the statement coverage in the passing tests.
  • Number of fixed hallucinated tests. We measure the number of tests that initially crash with “* is not a function” or “Reading property of undefined” errors and that subsequently De-Hallucinator turns into passing tests.
@@ 5.1.5. Hardware
We perform the experiments with the CodeGen v2.5 model and the GPT-3.5-turbo-0125 model on a machine equipped with two Nvidia T4 GPUs, each having 16GB of memory. The experiments with the UniXcoder, CodeGen, and the StarCoder+ models are performed on a machine with a single Nvidia Tesla V100 with 32GB of memory. Each machine has a 48-core Intel Xeon CPU clocked at 2.20GHz.
@@ 5.2. RQ1: Effectiveness of De-Hallucinator
Table 3. Effectiveness of De-Hallucinator in code completion compared to the baseline on four off-the-shelf LLMs. The bold numbers show statistically significant improvement over the initial prompt. The numbers in parentheses show the relative improvement over the baseline.
Prompt type UniXCoder 125M CodeGen v1 2B CodeGen v2.5 7B StarCoder+ 15B
Edit distance (lower is better):
Initial 52.4 40.0 47.2 44.6
RAG 46.8 (10.7%) 33.4 (16.5%) 31.6 (33.0%) 35.9 (19.4%)
Iterative 25.9 (50.6%) 30.7 (23.3%) 30.1 (36.3%) 33.5 (24.9%)
Edit similarity (lower is better):
Initial 33.4 43.6 43.9 33.2
RAG 37.3 (11.7%) 48.0 (10.0%) 49.4 (12.5%) 38.0 (14.3%)
Iterative 42.6 (27.5%) 48.9 (12.1%) 50.2 (14.2%) 39.7 (19.3%)
Exact API match (higher is better):
Initial 4.8 7.1 8.3 5.7
RAG 4.8 (0.0%) 10.2 (42.6%) 11.6 (39.1%) 7.5 (32.0%)
Iterative 5.9 (23.9%) 11.1 (55.3%) 13.4 (61.0%) 7.5 (32.0%)
Code Completion
@@
In the first set of experiments, we investigate to what extent De-Hallucinator improves code completions compared to the baseline. By default, we run De-Hallucinator with $k=3$ iterations and add $n=20$ API references into the prompt. As shown in Table 3, De-Hallucinator reduces the edit distance by 9.3 to 26.5 tokens, on average over the completions by the initial prompt, which is a relative improvement between 23.3% and 50.6% . This in turn translates to normalized edit similarity improvements of 12.1% to 27.5% relative to the baseline. Moreover, the approach relatively improves the exact API matches by 23.9% to 61.0%. For example, for the CodeGen v2.5 model, De-Hallucinator is able to predict 1.5 times more APIs correctly than the baseline. The approach shows statistically significant (using the Wilcoxon test and Pratt method) improvements over the baseline consistently for all metrics and all models. For example, Figure 5 shows a scenario where De-Hallucinator improves the completion. In this case the correct function is used by the model in the first try, but the order of parameters is wrong. By providing the API reference in the prompt, De-Hallucinator predicts the correct API usage, as shown in Fig. 5.
Table 4. Effectiveness of De-Hallucinator in test generation compared to TestPilot. The bold numbers show statistically significant improvement over the baseline. The numbers in parentheses show the relative improvement over the baseline.
Prompt type Passing tests Coverage Fixed hallucinations
Initial (TestPilot) 64.8 32.1 19.3
RAG & iterative 66.3 (3.1%) 33.7 (3.6%) 43.2 (94.0%)
Iterative 76.4 (17.9%) 37.0 (15.5%) 31.4 (63.2%)
Test Generation
In the second set of experiments, we evaluate the effectiveness of De-Hallucinator in test generation. We use $k=3$ and $n=3$ as default parameters for these experiments. Table 4 shows statistically significant (using Wilcoxon test and Pratt method) improvements for code coverage and fixed hallucinations by De-Hallucinator. Since RAG prompts use the initial prompt for retrieval, and because the initial prompt just contains a single signature, it is not as useful as iterative prompts. This is also reflected in Table 4 as lower coverage compared to only using iterative prompts. Note that TestPilot uses some prompt refiners, such as retrying with error message, and including usage snippets and function bodies. These refiners can mitigate some hallucinations, but not as much as De-Hallucinator, as shown in Section 5.2. For example, Figure 6 shows a test generated by TestPilot, which uses non-existing APIs, but after providing the API reference in the prompt De-Hallucinator generates a test using the correct API.
{minted}
python async def schedule_formatting(sources: Set[Path], fast: bool, write_back: WriteBack, mode: Mode, report: "Report", loop: asyncio.AbstractEventLoop, executor: "Executor") -> None: """Run formatting of ‘sources‘ in parallel using the provided ‘executor‘. (Use ProcessPoolExecutors for actual parallelism.) ‘write_back‘, ‘fast‘, and ‘mode‘ options are passed to :func:‘format_file_in_place‘. """ cache: Cache = if write_back not in (WriteBack.DIFF, WriteBack.COLOR_DIFF): cache = read_cache(mode) |sources, cached = filter_cached(sources, cache)| # <- baseline |sources, cached = filter_cached(cache, sources)| # <- ground truth {minted}python # API Reference: # filter_cached(cache: Cache, sources: Iterable[Path]) -> Tuple[Set[Path], Set[Path]] # Split an iterable of paths in ‘sources‘ into two sets. The first contains paths of files that modifi … async def schedule_formatting(sources: Set[Path], fast: bool, write_back: WriteBack, mode: Mode, report: "Report", loop: asyncio.AbstractEventLoop, executor: "Executor") -> None: """Run formatting of ‘sources‘ in parallel using the provided ‘executor‘. (Use ProcessPoolExecutors for actual parallelism.) ‘write_back‘, ‘fast‘, and ‘mode‘ options are passed to :func:‘format_file_in_place‘. """ cache: Cache = if write_back not in (WriteBack.DIFF, WriteBack.COLOR_DIFF): cache = read_cache(mode) |sources, cached = filter_cached(cache, sources)| #<-|De-Hallucinator|
@@
{minted}
[highlightcolor=bgred,highlightlines=11,14,17,19]javascript let mocha = require(’mocha’); let assert = require(’assert’); let js_sdsl = require(’js-sdsl’); // class js-sdsl.Deque(t = [], i = 1 << 12) describe(’test js_sdsl’, function() it(’test js-sdsl.Deque’, function(done) let deque = new js_sdsl.Deque([1, 2, 3, 4, 5]); assert.equal(deque.size(), 5); assert.equal(deque.front(), 1); assert.equal(deque.back(), 5); deque.push_back(6); assert.equal(deque.size(), 6); assert.equal(deque.back(), 6); deque.push_front(0); assert.equal(deque.size(), 7); assert.equal(deque.front(), 0); assert.equal(deque.pop_back(), 6); assert.equal(deque.size(), 6); assert.equal(deque.pop_front(), 0); assert.equal(deque.size(), 5); done(); ) ) {minted}[highlightcolor=bggreen,highlightlines=13]javascript let mocha = require(’mocha’); let assert = require(’assert’); let js_sdsl = require(’js-sdsl’); // class js-sdsl.Deque(t = [], i = 1 << 12) // API Reference: // js-sdsl.Deque.prototype.pushFront(t) // js-sdsl.Deque.prototype.front() // js-sdsl.Queue.prototype.front() describe(’test js_sdsl’, function() it(’test js-sdsl.Deque’, function(done) let deque = new js_sdsl.Deque([1, 2, 3]); assert.equal(deque.front(), 1); deque.pushFront(0); assert.equal(deque.front(), 0); done(); ) )
Figure 6. Test generated by TestPilot using GPT-3.5-turbo (top), and the iterative prompt by De-Hallucinator resulting in correct usage of APIs (bottom).
@@ 5.3. RQ2: Correct Retrieval of API References
Code Completion
To better understand the effectiveness of De-Hallucinator, we investigate how often the approach successfully augments the prompt with the correct API references. Answering this question for all code completion tasks and all LLMs is difficult, because comparing the API references to the API usages is non-trivial due to different ways of importing APIs and passing arguments. Instead, for code completion, we manually inspect a sample of 20 completion tasks per LLM and count the number of times an API used in the ground truth is successfully added to the prompt by De-Hallucinator.
@@
Missing Expected API ref. added
/wrong RAG Iter. 1 Iter. 2 Iter. 3
@@
In our inspected samples for code completion, there are in total between 22 and 25 API usages (a completion can contain multiple API usages). The new prompt generated by our approach contains the correct API between two and six times, as shown in Table 5. Moreover, for CodeGen and CodeGen v2.5 there are five completion tasks where the completion from the initial prompt either misses the APIs or uses them incorrectly, but in the completions from RAG or iterative prompts, the API reference section of the prompt contains the correct API. The same happens for UniXcoder in four tasks and for StarCoder+ in two tasks. For cases where the approach fails to add the correct API reference into the prompt, the main reason is that the initial completion has low relevance w.r.t. the ground truth.
Test Generation
Since for the test generation task there are no ground truths for the generated code pieces, manual inspection is infeasible. Instead, we measure the number of passing tests and the number of non-crashing failing tests from an iterative prompt, as a proxy of success for the retrieval. Note that an iterative prompt is only created when an initial prompt causes a crash with one of the specified errors in Section 4. We observe that from the 622 instances where an iterative prompt is generated, 16.7% of iterative prompts result in a passing test, and 26.8% of iterative prompts result in a test without hallucinations. These results are also in line with the manual inspections of the code completion task above.
Overall, the results show that De-Hallucinator is able to successfully augment the prompt with the correct API references in many cases.
@@ 5.4. RQ3: Impact of the Hyperparameters
This research question evaluates the impact of the two main parameters of our approach. First, we consider the number $k$ of iterations of iterative prompting. For both code completion and test generation, we run the approach with $k\in\{1,2,3\}$. As shown in Fig. 7, the first iteration provides significant improvement over the baseline, but the gain is reduced with further iterations. Higher values of $k$ are beneficial when the model cannot immediately predict a relevant code, but upon presenting the first round of API references, the model responds with more relevant outputs. At the same time, even $k=1$ provides clear improvements over the baseline, which makes De-Hallucinator useful even in scenarios where the cost of querying the model is high. We choose $k=3$ for all other experiments in this paper.
Figure 7. Effects of $k$ and $n$ for code completion.
Figure 8. Effects of $k$ and $n$ for test generation.
Second, we study the impact of the maximum number $n$ of API references that we add to the prompt. Depending on the task, the optimal number of API references in the prompt varies. Setting low values for $n$ can result in missing relevant context, whereas adding many API references can confuse the model, while also costing context space. We perform experiments with $n\in\{2,10,20,40\}$ for code completion, and $n\in\{3,5,10\}$ for test generation. Figure 7 shows the results of code completion, with exact API match peaking between $n=10$ and $n=20$. As a default in the rest of the paper, we use $n=20$ for code completion. Figure 8 shows the effects of $n$ on coverage of the generated tests. The peak in this case is at $n=3$, which we use as the parameter for other experiments in this paper.